OpenStudy (anonymous):

Can anyone explain the Fibonacci sequence, I understand that the third term is suppose to be the sum of the two previous numbers, but this is becoming confusing. The first four terms in a pattern are 4, 7, 10, and 13. What is the fifth term in the pattern? Shouldn't it be; 4,7,11,18?

OpenStudy (anonymous):

@iGreen

OpenStudy (amistre64):

f1 = 4 f2 = 7 f3 = f2 + f1 f4 = f3 + f2 f5 = f4 + f3

OpenStudy (amistre64):

your pattern is not fibby, why are you wanting to make it a fibby?

OpenStudy (anonymous):

Ohh, because, I was learning about it in the section. It never occurred to me that it wouldn't be. . .

OpenStudy (amistre64):

it looks arithmetic to me ....

OpenStudy (amistre64):

writing out the difference tiers tends to be useful to me when trying to find a pattern 4 7 10 13 3 3 3 looks to be adding 3 each time

OpenStudy (anonymous):

Okay, I see the pattern now that I've stopped applying the Fibonacci technique, thanks.

OpenStudy (amistre64):

:)